# ESPFormer: Doubly-Stochastic Attention with Expected Sliced Transport Plans

### v1

@@ Abstract
While self-attention has been instrumental in the success of Transformers, it can lead to over-concentration on a few tokens during training, resulting in suboptimal information flow. Enforcing doubly-stochastic constraints in attention matrices has been shown to improve structure and balance in attention distributions. However, existing methods rely on iterative Sinkhorn normalization, which is computationally costly. In this paper, we introduce a novel, fully parallelizable doubly-stochastic attention mechanism based on sliced optimal transport, leveraging Expected Sliced Transport Plans (ESP). Unlike prior approaches, our method enforces double stochasticity without iterative Sinkhorn normalization, significantly enhancing efficiency. To ensure differentiability, we incorporate a temperature-based soft sorting technique, enabling seamless integration into deep learning models. Experiments across multiple benchmark datasets, including image classification, point cloud classification, sentiment analysis, and neural machine translation, demonstrate that our enhanced attention regularization consistently improves performance across diverse applications.
@@ 1 Introduction
The debut of Transformers [51] marked a turning point in artificial intelligence and machine learning. Self-attention mechanisms excel at modeling the interactions among features, allowing Transformers to generate highly expressive and context-rich representations that accurately capture the essence of input data [18]. Although first developed to perform Natural Language Processing (NLP) tasks, the Transformer architecture has been adapted to a wide range of domains such as computer vision [7, 48, 49, 50], graphs [57, 42, 45], point clouds [61, 40], and biological sequences [15, 1, 43, 26, 12]. Over the years, a considerable amount of research has been devoted to enhancing the classic Transformer, focusing, among others, on better positional encoding [8, 56], efficiency of the attention mechanisms [52, 3], and variants of self-attention [14, 13].
@@
Self-attention produces a row-stochastic matrix, which can lead to a few tokens dominating the attention distribution. To mitigate this, enforcing doubly-stochastic attention ensures a more balanced distribution across tokens. To that end, Sander et al. [44] introduced Sinkformer, replacing the softmax normalization in the classic Transformer with Sinkhorn’s algorithm [46], resulting in a doubly-stochastic attention matrix. Sander et al. [44] establish a connection between self-attention matrices and the optimal transport problem by theoretically demonstrating that Sinkformers can be interpreted as a Wasserstein gradient flow for an energy minimization in the infinite depth limit. In fact, there exists a fundamental compatibility between transport plans and Transformer architectures, rooted in their shared property of permutation-equivariance. In particular, due to their inherent doubly-stochastic nature, transport plans are good candidates for doubly-stochastic attention mechanisms.
@@
Nevertheless, computing optimal transport plans between keys and queries is computationally expensive, with a complexity of $\mathcal{O}(N^{3})$ for $N$ tokens. A more scalable approach is entropy-regularized transport [4], which leverages the Sinkhorn algorithm to iteratively approximate the transport plan, obtaining a complexity $\mathcal{O}(SN^{2})$, where $S$ denotes the number of iterations. The Sinkhorn algorithm [46], as used in Sinkformers [44], repeatedly applies a series of row and column normalization steps until it reaches a desired level of convergence. Such an approach may introduce computational inefficiencies in scenarios where a significant number of iterations are needed for the normalization to converge.
@@
A more computationally efficient alternative is the calculation of optimal transport plans for one-dimensional distributions, with a complexity of $\mathcal{O}(N\log N)$. This has motivated a large body of work on sliced optimal transport methods that compare distributions by comparing their slices, i.e., one-dimensional marginals [41, 20, 21, 22, 6, 35, 34]. Although sliced optimal transport methods offer efficient metrics between distributions, they do not explicitly construct transport plans. Recent studies have addressed this limitation by developing methods to construct transport plans through slicing [30, 27]. Liu et al. [27] introduce the Expected Sliced Transport Plan (ESP), which leverages slicing to define a computationally efficient metric while crucially providing an explicit transport plan by aggregating lifted plans from all slices. This makes ESP a strong candidate for doubly-stochastic attention mechanisms.
@@
  • We demonstrate the compatibility of our proposed attention mechanism with the recently introduced differential attention architecture [55].
@@ 2 Background and Related Work
In this section, we first provide an overview of the existing variants of softmax attention, including the doubly-stochastic attention using Sinkhorn’s algorithm [44]. Then, we shift our focus to the fundamentals of sliced optimal transport, with soft sorting reviewed for algorithmic concerns. Finally, we provide an overview of Expected Sliced Transport Plans [27] and soft sorting [38], which serve as the cornerstones of our proposed doubly-stochastic attention mechanism, ESPFormer.
@@ 2.1 Variants of Softmax Attention
At the heart of the Transformer architecture is the self-attention operation, a crucial component that enables dynamic pairwise interactions among tokens. In essence, it allows each position to ‘attend” to all others, with the degree of attention determined by how similar their representations are. Formally, let $W_{Q},W_{K}\in\mathbb{R}^{m\times d},W_{V}\in\mathbb{R}^{d\times d}$ denote the query, key, and value matrices, respectively. Then, for a sequence $(x_{1},x_{2},\cdots,x_{N}),x_{i}\in\mathbb{R}^{d},\forall i$, the output of the attention function for the $i$-th row, $x_{i}$, can be written as
@@
where $\text{sim}(\cdot,\cdot)$ can be any similarity function, and the normalization for the similarities is applied row-wise. The classic self-attention mechanism [51] leverages the softmax function to perform this row-wise normalization, i.e., the softmax of attention matrix $C$ with $C_{i,j}=(W_{Q}x_{i})^{T}W_{K}x_{j}$ can be interpreted as row-wise normalization of $\exp(C)$. Alternative normalization operators have also been proposed in the literature. Some focus on normalizations that share the same properties as softmax but produce sparse outputs, such as SparseMax [31] and SparseK [28]. SparseMax seeks the Euclidean projection of the input into the probabilistic simplex. Since this projection tends to hit the boundary of the simplex, SparseMax will output sparse probabilities. SparseK extends SparseMax by replacing the probabilistic simplex with a $k$-sum constraint. Others use different $\text{sim}(\cdot,\cdot)$ functions in (1) to achieve linear complexity: Katharopoulos et al. [17] and Han et al. [11] use kernel smoothers to linearize the similarity calculations, and Li et al. [25] propose to approximate the exponential function in softmax by the first-order Taylor expansion.
Another line of research approaches the attention from the alignment/matching perspective and specifically utilizes concepts from the Optimal Transport (OT) theory. Zhang et al. [58] propose to enforce the alignment between the distributions of keys and queries by adding a penalty term in the training objective, where the Jensen–Shannon (JS) divergence, the Wasserstein distance, and bi-directional conditional transport [62] are considered to define the matching cost. Xu and Chen [54] introduce a multimodal co-attention method that relies on OT to match patches of images and gene embeddings for the survival prediction task. Zhang et al. [60] reinterpret slot attention within the OT framework and propose to enhance the performance of slot attention by minimizing the entropy of the Sinkhorn divergence between two multisets, one containing inputs and the other containing slot features.
@@ 2.2 Doubly-Stochastic Attention
In their pioneering work, Sander et al. [44] empirically observed that, during training, the row-stochastic attention matrices in classical Transformers tend to become approximately doubly-stochastic, with most column sums approaching 1. This suggests that Transformers inherently learn to distribute attention more evenly across tokens. In light of this finding, the authors propose the Sinkformer architecture, which replaces the softmax operation by the Sinkhorn’s algorithm [46, 4, 37] to enforce doubly-stochastic attention as an informative prior. They show that although both classic Transformers and Sinkformers can be understood as models that operate on discrete probability distributions, the Sinkformers have a special property that, in the infinite depth limit, they behave as a Wasserstein gradient flow for an energy minimization.
@@ 2.3 Sliced Optimal Transport
where $\Gamma(\mu^{1},\mu^{2})$ denotes the set of all couplings between $\mu^{1}$ and $\mu^{2}$, and $c:\Omega^{2}\rightarrow\mathbb{R}_{+}$ is a lower semi-continuous function. Specifically, when $c(\cdot,\cdot)$ is the $p$-th power of a metric, the $p$-Wasserstein distance is defined as
@@
In practical settings, the Wasserstein distance between discrete probability measures can be obtained by solving a linear program [37] with an expensive computational complexity of $\mathcal{O}(N^{3}\log N)$ for sample size $N$. Therefore, faster alternatives have been extensively studied, including entropy-regularized optimal transport [4] and sliced optimal transport [20, 21, 6, 22, 35, 34]. Sinkhorn’s algorithm at the core of Sinkformer is used to solve the entropy-regularized optimal transport problem with an improved complexity of $\mathcal{O}(N^{2}\log N)$.
Sliced optimal transport operates by projecting high-dimensional distributions onto one-dimensional slices, leveraging the key property that, in the one-dimensional case, a unique optimal transport plan exists, and the $p$-Wasserstein distance has a closed-form solution [41],
$W_{p}(\mu^{1},\mu^{2})=\int_{0}^{1}\left|F^{-1}_{\mu^{1}}(u)-F^{-1}_{\mu^{2}}(u)\right|^{p}\text{d}u,$ (4)
where $F^{-1}_{\mu^{1}}$ and $F^{-1}_{\mu^{2}}$ are the quantile functions of $\mu^{1}$ and $\mu^{2}$ respectively, and the optimal transport map is given by $T(x)=F^{-1}_{\mu^{2}}(F_{\mu^{1}})$ when $\mu^{1}$ is continuous. For empirical measures with $N$ samples, the quantile functions correspond to the sorted samples that can be calculated in $\mathcal{O}(N\log N)$. Then, by integrating the 1-dimensional Wasserstein distance over a set of $L$ slices, the sliced Wasserstein distance reduces the computational cost significantly to $\mathcal{O}(LN\log N)$.
@@ 2.4 Expected Sliced Transport Plans
Liu et al. [27] recently proposed the Expected Sliced Transport Plan (ESP), which defines an efficient transport plan as an aggregation of lifted optimal transport plans on 1-dimensional slices. Let $\mu^{1}=\sum_{x\in\mathbb{R}^{d}}p(x)\delta_{x}$ be a discrete probability measure in $\mathcal{P}(\mathbb{R}^{d})$, that is, $p(x)\geq 0$ for all $x\in\mathbb{R}^{d}$ and $\sum_{x\in\mathbb{R}^{d}}p(x)=1$. We further assume that $p(x)\neq 0$ for at most countable many points $x\in\mathbb{R}^{d}$. Similarly, let $\mu^{2}=\sum_{y\in\mathbb{R}^{d}}q(y)\delta_{y}$ be another probability measure in $\mathcal{P}(\mathbb{R}^{d})$ with at most countable support. For a given $\theta\in\mathbb{S}^{d-1}$, the projected measures $\theta_{\#}\mu^{1}$ and $\theta_{\#}\mu^{2}$ are 1-dimensional probability measures in $\mathcal{P}(\mathbb{R})$, and there exists a unique optimal transport plan $\Lambda_{\theta}^{\mu^{1},\mu^{2}}$ between them. Equivalently, $\theta_{\#}\mu^{1}$ and $\theta_{\#}\mu^{2}$ can be interpreted as probability measures over a quotient space $\mathbb{R}^{d}/\sim_{\theta}$, where $\sim_{\theta}$ is defined as follows:
@@
as each point on the slice $\mathbb{R}$ corresponds to an equivalent class of points in $\mathbb{R}^{d}$ that gets mapped to it by $\theta$. With a slight abuse of notation, we denote the equivalent class of $x\in\mathbb{R}^{d}$ by $\bar{x}^{\theta}$, referring to either a point in the quotient space $\mathbb{R}^{d}/\sim_{\theta}$ or the set of points $\{x^{\prime}\in\mathbb{R}^{d}:\theta\cdot x^{\prime}=\theta\cdot x\}$ interchangeably. Then, we can write $\theta_{\#}\mu^{1}=\sum_{\bar{x}^{\theta}\in\mathbb{R}/\sim_{\theta}}P(\bar{x}^{\theta})\delta_{\bar{x}^{\theta}}$, where $P(\bar{x}^{\theta})=\sum_{x^{\prime}\in\bar{x}^{\theta}}p(x^{\prime})$, and $\theta_{\#}\mu^{2}=\sum_{\bar{y}^{\theta}\in\mathbb{R}/\sim_{\theta}}Q(\bar{y}^{\theta})\delta_{\bar{y}^{\theta}}$, where $Q(\bar{y}^{\theta})=\sum_{y^{\prime}\in\bar{y}^{\theta}}q(y^{\prime})$.
This quotient space interpretation of the 1-dimensional distributions $\theta_{\#}\mu^{1}$ and $\theta_{\#}\mu^{2}$ allows us to construct a lifted transport plan in the original space $\mathbb{R}^{d}$ using the optimal transport plan $\Lambda_{\theta}^{\mu^{1},\mu^{2}}$,
@@
Liu et al. [27] have shown that the associated cost,

### v2

@@ Abstract
While self-attention has been instrumental in the success of Transformers, it can lead to over-concentration on a few tokens during training, resulting in suboptimal information flow. Enforcing doubly-stochastic constraints in attention matrices has been shown to improve structure and balance in attention distributions. However, existing methods rely on iterative Sinkhorn normalization, which is computationally costly. In this paper, we introduce a novel, fully parallelizable doubly-stochastic attention mechanism based on sliced optimal transport, leveraging Expected Sliced Transport Plans (ESP). Unlike prior approaches, our method enforces doubly stochasticity without iterative Sinkhorn normalization, significantly enhancing efficiency. To ensure differentiability, we incorporate a temperature-based soft sorting technique, enabling seamless integration into deep learning models. Experiments across multiple benchmark datasets, including image classification, point cloud classification, sentiment analysis, and neural machine translation, demonstrate that our enhanced attention regularization consistently improves performance across diverse applications. Our implementation code can be found at https://github.com/dariansal/ESPFormer.
Machine Learning, ICML
[…]
@@ 1 Introduction
The debut of Transformers (Vaswani et al., 2017) marked a turning point in artificial intelligence and machine learning. Self-attention mechanisms excel at modeling the interactions among features, allowing Transformers to generate highly expressive and context-rich representations that accurately capture the essence of input data (Khan et al., 2022). Although first developed to perform Natural Language Processing (NLP) tasks, the Transformer architecture has been adapted to a wide range of domains such as computer vision (Dosovitskiy et al., 2021; Touvron et al., 2021a, b, 2022), graphs (Yun et al., 2019; Rampášek et al., 2022; Shirzad et al., 2023), point clouds (Zhao et al., 2021; Qin et al., 2022), and biological sequences (Jumper et al., 2021; Abramson et al., 2024; Rives et al., 2021; Lin et al., 2023; Hayes et al., 2025). Over the years, a considerable amount of research has been devoted to enhancing the classic Transformer, focusing on better positional encoding (Dwivedi et al., 2021; Ying et al., 2021), efficiency of the attention mechanisms (Wang et al., 2020; Choromanski et al., 2020), and variants of self-attention (Hou et al., 2019; Ho et al., 2019), among others.
@@
Self-attention produces a row-stochastic matrix, which can lead to a few tokens dominating the attention distribution. To mitigate this, enforcing doubly-stochastic attention ensures a more balanced distribution across tokens. To that end, Sander et al. (2022) introduced Sinkformer, replacing the softmax normalization in the classic Transformer with Sinkhorn’s algorithm (Sinkhorn, 1964), resulting in a doubly-stochastic attention matrix. Sander et al. (2022) establish a connection between self-attention matrices and the optimal transport problem by theoretically demonstrating that Sinkformers can be interpreted as a Wasserstein gradient flow for an energy minimization in the infinite depth limit. Indeed, doubly stochastic attention can be interpreted as a transport plan between queries and keys, further strengthening the connection between optimal transport and attention mechanisms. Owing to their inherent doubly stochastic structure, transport plans (i.e., couplings between keys and queries) naturally serve as strong candidates for attention matrices in this setting.
@@
Nevertheless, computing optimal transport plans between keys and queries is computationally expensive, with a general complexity of $\mathcal{O}(N^{3})$ for $N$ tokens. A more scalable approach is entropy-regularized transport (Cuturi, 2013), which leverages the Sinkhorn algorithm to iteratively approximate the transport plan, obtaining a complexity $\mathcal{O}(SN^{2})$, where $S$ denotes the number of iterations. The Sinkhorn algorithm (Sinkhorn, 1964), as used in Sinkformers (Sander et al., 2022), repeatedly applies a series of row and column normalization steps until it reaches a desired level of convergence. Such an approach may introduce computational inefficiencies in scenarios where a significant number of iterations are needed for the normalization to converge.
@@
A more computationally efficient alternative is the calculation of optimal transport plans for one-dimensional distributions, with a complexity of $\mathcal{O}(N\log N)$. This has motivated a large body of work on sliced optimal transport methods that compare distributions by comparing their slices, i.e., one-dimensional marginals (Rabin et al., 2012; Kolouri et al., 2016, 2018, 2019; Deshpande et al., 2019; Nguyen et al., 2022; Nguyen & Ho, 2024). Although sliced optimal transport methods offer efficient metrics between distributions, they do not explicitly construct transport plans. Recent studies have addressed this limitation by developing methods to construct transport plans through slicing (Mahey et al., 2024; Liu et al., 2025). Liu et al. (2025) introduce the Expected Sliced Transport Plan (ESP), which leverages slicing to define a computationally efficient metric while crucially providing an explicit transport plan by aggregating lifted plans from all slices. This makes ESP a strong candidate for doubly-stochastic attention mechanisms.
@@
  • We show that by fine‐tuning pretrained models with an exponential temperature annealing schedule and then switching to hard sorting at inference, we obtain exact doubly stochastic matrices reducing complexity from $O(N^{2})$ to $O(LN\log N)$, and achieve consistent accuracy gains on the Cats vs. Dogs dataset.
  • We demonstrate the compatibility of our proposed attention mechanism with the recently introduced differential attention architecture (Ye et al., 2025).
@@ 2 Background and Related Work
In this section, we first provide an overview of the existing variants of softmax attention, including the doubly-stochastic attention using Sinkhorn’s algorithm (Sander et al., 2022). Then, we shift our focus to the fundamentals of sliced optimal transport, with soft sorting reviewed for algorithmic concerns. Finally, we provide an overview of Expected Sliced Transport Plans (Liu et al., 2025) and soft sorting (Prillo & Eisenschlos, 2020), which serve as the cornerstones of our proposed doubly-stochastic attention mechanism, ESPFormer.
@@ 2.1 Variants of Softmax Attention
At the heart of the Transformer architecture is the self-attention operation, a crucial component that enables dynamic pairwise interactions among tokens. In essence, it allows each position to “attend” to all others, with the degree of attention determined by how similar their representations are. Formally, let $W_{Q},W_{K}\in\mathbb{R}^{m\times d},W_{V}\in\mathbb{R}^{d\times d}$ denote the query, key, and value matrices, respectively. Then, for a sequence $(x_{1},x_{2},\cdots,x_{N}),x_{i}\in\mathbb{R}^{d},\forall i$, the output of the attention function for the $i$-th row, $x_{i}$, can be written as
@@
where $\text{sim}(\cdot,\cdot)$ can be any similarity function, and the normalization for the similarities is applied row-wise. The classic self-attention mechanism (Vaswani et al., 2017) leverages the softmax function to perform this row-wise normalization, i.e., the softmax of attention matrix $C$ with $C_{i,j}=(W_{Q}x_{i})^{T}W_{K}x_{j}$ can be interpreted as row-wise normalization of $\exp(C)$. Alternative normalization operators have also been proposed in the literature. Some focus on normalizations that share the same properties as softmax but produce sparse outputs, such as SparseMax (Martins & Astudillo, 2016) and SparseK (Lou et al., 2024). SparseMax seeks the Euclidean projection of the input into the probabilistic simplex. Since this projection tends to hit the boundary of the simplex, SparseMax will output sparse probabilities. SparseK extends SparseMax by replacing the probabilistic simplex with a $k$-sum constraint. Others use different $\text{sim}(\cdot,\cdot)$ functions in (1) to achieve linear complexity: Katharopoulos et al. (2020) and Han et al. (2023) use kernel smoothers to linearize the similarity calculations, and Li et al. (2020) propose to approximate the exponential function in softmax by the first-order Taylor expansion.
Another line of research approaches the attention from the alignment/matching perspective and specifically utilizes concepts from the Optimal Transport (OT) theory. Zhang et al. (2021) propose to enforce the alignment between the distributions of keys and queries by adding a penalty term in the training objective, where the Jensen–Shannon (JS) divergence, the Wasserstein distance, and bi-directional conditional transport (Zheng & Zhou, 2021) are considered to define the matching cost. Xu & Chen (2023) introduce a multimodal co-attention method that relies on OT to match patches of images and gene embeddings for the survival prediction task. Zhang et al. (2023) reinterpret slot attention within the OT framework and propose to enhance the performance of slot attention by minimizing the entropy of the Sinkhorn divergence between two multisets, one containing inputs and the other containing slot features.
@@ 2.2 Doubly-Stochastic Attention
In their pioneering work, Sander et al. (2022) empirically observed that, during training, the row-stochastic attention matrices in classical Transformers tend to become approximately doubly-stochastic, with most column sums approaching 1. This suggests that Transformers inherently learn to distribute attention more evenly across tokens. In light of this finding, the authors propose the Sinkformer architecture, which replaces the softmax operation by the Sinkhorn’s algorithm (Sinkhorn, 1964; Cuturi, 2013; Peyré et al., 2019) to enforce doubly-stochastic attention as an informative prior. They show that although both classic Transformers and Sinkformers can be understood as models that operate on discrete probability distributions, the Sinkformers have a special property that, in the infinite depth limit, they behave as a Wasserstein gradient flow for an energy minimization.
@@ 2.3 Sliced Optimal Transport
where $\Gamma(\mu^{1},\mu^{2})$ denotes the set of all couplings between $\mu^{1}$ and $\mu^{2}$, and $c:\Omega^{2}\rightarrow\mathbb{R}_{+}$ is a lower semi-continuous function representing the transportation cost. Specifically, when $c(\cdot,\cdot)$ is the $p$-th power of a metric, the $p$-Wasserstein distance is defined as
@@
In practical settings, the Wasserstein distance between discrete probability measures can be obtained by solving a linear program (Peyré et al., 2019) with an expensive computational complexity of $\mathcal{O}(N^{3}\log N)$ for sample size $N$. Therefore, faster alternatives have been extensively studied, including entropy-regularized optimal transport (Cuturi, 2013) and sliced optimal transport (Rabin et al., 2012; Kolouri et al., 2016, 2018; Deshpande et al., 2019; Kolouri et al., 2019; Nguyen et al., 2022; Nguyen & Ho, 2024). Sinkhorn’s algorithm at the core of Sinkformer is used to solve the entropy-regularized optimal transport problem with an improved complexity of $\mathcal{O}(N^{2}\log N)$.
Sliced optimal transport operates by projecting high-dimensional distributions onto one-dimensional slices, leveraging the key property that, in the one-dimensional case, a unique optimal transport plan exists, and the $p$-Wasserstein distance has a closed-form solution (Rabin et al., 2012),
$W_{p}(\mu^{1},\mu^{2})=\left(\int_{0}^{1}\left|F^{-1}_{\mu^{1}}(u)-F^{-1}_{\mu^{2}}(u)\right|^{p}\text{d}u\right)^{\frac{1}{p}},$ (4)
where $F^{-1}_{\mu^{1}}$ and $F^{-1}_{\mu^{2}}$ are the quantile functions of $\mu^{1}$ and $\mu^{2}$ respectively, and the optimal transport map is given by $T(x)=F^{-1}_{\mu^{2}}(F_{\mu^{1}}(x))$ when $\mu^{1}$ is absolutely continuous. For empirical measures with $N$ samples, the quantile functions correspond to the sorted samples that can be calculated in $\mathcal{O}(N\log N)$. Then, by integrating the 1-dimensional Wasserstein distance over a set of $L$ slices, the sliced Wasserstein distance reduces the computational cost significantly to $\mathcal{O}(LN\log N)$.
@@ 2.4 Expected Sliced Transport Plans
Liu et al. (2025) recently proposed the Expected Sliced Transport Plan (ESP), which defines an efficient transport plan as an aggregation of lifted optimal transport plans on 1-dimensional slices. Let $\mu^{1}=\sum_{x\in\mathbb{R}^{d}}p(x)\delta_{x}$ be a discrete probability measure in $\mathcal{P}(\mathbb{R}^{d})$, that is, $p(x)\geq 0$ for all $x\in\mathbb{R}^{d}$ and $\sum_{x\in\mathbb{R}^{d}}p(x)=1$. We further assume that $p(x)\neq 0$ for at most countable many points $x\in\mathbb{R}^{d}$. Similarly, let $\mu^{2}=\sum_{y\in\mathbb{R}^{d}}q(y)\delta_{y}$ be another probability measure in $\mathcal{P}(\mathbb{R}^{d})$ with at most countable support. For a given $\theta\in\mathbb{S}^{d-1}$, the projected measures $\theta_{\#}\mu^{1}$ and $\theta_{\#}\mu^{2}$ are 1-dimensional probability measures in $\mathcal{P}(\mathbb{R})$, and there exists a unique optimal transport plan $\Lambda_{\theta}^{\mu^{1},\mu^{2}}$ between them. Equivalently, $\theta_{\#}\mu^{1}$ and $\theta_{\#}\mu^{2}$ can be interpreted as probability measures over a quotient space $\mathbb{R}^{d}/\sim_{\theta}$, where $\sim_{\theta}$ is defined as follows:
@@
as each point on the slice $\mathbb{R}$ corresponds to an equivalent class of points in $\mathbb{R}^{d}$ that gets mapped to it by $\theta$. With a slight abuse of notation, we denote the equivalent class of $x\in\mathbb{R}^{d}$ by $\bar{x}^{\theta}$, referring to either a point in the quotient space $\mathbb{R}^{d}/\sim_{\theta}$ or the set of points $\{x^{\prime}\in\mathbb{R}^{d}:\theta\cdot x^{\prime}=\theta\cdot x\}$ interchangeably. Then, we can write $\theta_{\#}\mu^{1}=\sum_{\bar{x}^{\theta}\in\mathbb{R}/\sim_{\theta}}P(\bar{x}^{\theta})\delta_{\bar{x}^{\theta}}$, where $P(\bar{x}^{\theta})=\sum_{x^{\prime}\in\bar{x}^{\theta}}p(x^{\prime})$, and $\theta_{\#}\mu^{2}=\sum_{\bar{y}^{\theta}\in\mathbb{R}/\sim_{\theta}}Q(\bar{y}^{\theta})\delta_{\bar{y}^{\theta}}$, where $Q(\bar{y}^{\theta})=\sum_{y^{\prime}\in\bar{y}^{\theta}}q(y^{\prime})$. This quotient space interpretation of the 1-dimensional distributions $\theta_{\#}\mu^{1}$ and $\theta_{\#}\mu^{2}$ allows us to construct a lifted transport plan in the original space $\mathbb{R}^{d}$ using the optimal transport plan $\Lambda_{\theta}^{\mu^{1},\mu^{2}}$,
@@
Liu et al. (2025) have shown that the associated cost,